OpenStudy (anonymous):

wo years ago Adam was 6 times as old as his son. In 3 years he will be 3 1/2 times as old as his son. Find their present ages.

OpenStudy (anonymous):

How to solve this using 2 variables?

OpenStudy (anonymous):

what should i do next?

OpenStudy (perl):

Question: **two years ago Adam was 6 times as old as his son. In 3 years he will be 3 1/2 times as old as his son. Find their present ages. Assume age of adam now = x adam's son age now = y. x-2 = 6(y-2) x+3 = 3.5(y+3)

OpenStudy (perl):

sorry i made a mistake earlier, we have subtract 2 from x and y, and add 3 to both x and y (use parentheses)

OpenStudy (perl):

because it says 2 years ago, both adam and his son's age decreases by 2

OpenStudy (anonymous):

so then it becomes x-2=6y-12 and x+3=3.5y+10.5 ?

OpenStudy (perl):

yes

OpenStudy (perl):

now you can solve by elimination (subtract the two equations and eliminate x), or you can solve for x in both equations

OpenStudy (perl):

x = 6y -10 x = 3.5y + 7.5 so they both equal to x (and so equal to each other) 6y-10 = 3.5y + 7.5 2.5y = 17.5 y = 7 x = 6*7-10 = 32

OpenStudy (perl):

this problem is a good example of algebraic thinking

OpenStudy (anonymous):

ok. thanks!